OpenStudy (anonymous):

H3P4: Diode Limiter is there any one to guide me to solve it?

OpenStudy (anonymous):

When voltage source is +7.5 then only D1 will work,now apply thevenin's process and find the current.and for -7.5 only D2 will work.

OpenStudy (anonymous):

For D2=(-V2/R)-((V2+Vin)/R) Remember who is conducting and use LCK

OpenStudy (anonymous):

v2=3 r=5600 vin=???? 7.5/ \[\sqrt{3}\]

OpenStudy (anonymous):

Vin = 7.5

OpenStudy (anonymous):

ID2=-0.002411A its wrong answer

OpenStudy (anonymous):

(-Vs=V2)/R-(-V2)R = (-7.5+3)/5600 -(-3)/5600 = -0.000268 however the diode is going in the opposite direction therefore the ans should be 0.000268

OpenStudy (anonymous):

take the negative sign and thats it

OpenStudy (anonymous):

Anyone can help me with the lab?